# Understanding the approach to thermalization in a non-Abelian gauge theory

### v1

@@ Abstract
We measure the maximal Lyapunov exponent $\lambda_{L}$ of physical states in a SU(2) gauge theory both in and out-of-thermal equilibrium conditions using ab-initio lattice techniques. A non-equilibrium state is realized starting from an over-occupied initial condition for low momentum, soft gluons whereas the thermal state comprises of strongly interacting soft gluons at temperatures where these are well separated from the hard momentum modes. A spectra of positive Lyapunov exponents is observed in both these states, similar to a chaotic dynamical system. From the Kolmogorov-Sinai entropy rate measured in terms of this spectrum, we estimate the typical time-scale of $\sim 0.50(3)$ fm/c to achieve thermalization at $T\sim 600$ MeV starting from the non-thermal state. We also measure, for the first time, the $\lambda_{L}$ for long wavelength critical modes of SU(2) using the out-of-time-ordered correlation functions of a classical $Z_{2}$ scalar field theory, which shares the same universal behavior near the deconfinement phase transition. The $\lambda_{L}$ is observed to maximize at the transition temperature.
@@
Non-equilibrium quantum field theory , QCD , non-Abelian gauge theory , chaos , out-of-time-ordered correlation
@@ 1 Motivation
Understanding the mechanism of thermalization in strongly interacting systems is a challenging research problem in theoretical physics [1, 2] as it would be able to explain physics at different length scales, e.g. in heavy-ion collisions, cold atom traps or in the early universe. Starting from a non-equilibrium state, a necessary condition for eventually reaching a thermal state is a mixing between different regions of the phase space. A mixing in the phase space can occur if there is a mechanical instability in the system [3], which is characterized by at least one positive Lyapunov exponent. Furthermore in the thermal state the entropy or the information content [4] is maximized [5]. Except for a few simple systems [6, 7] it is not very well understood how the entropy evolves as a function of time, starting from a chaotic non-equilibrium state and eventually maximizing itself in a thermal state. One important observable is the rate of change of entropy, known as the Kolmogorov-Sinai (KS) entropy rate [8], which equals the sum of all positive Lyapunov exponent for Hamiltonian systems [9]. This is known to be true for dissipative systems that has an attractor solution. Hence calculating the Lyapunov exponents in chaotic dynamical systems is crucial to understand if, and in what time-scales these would eventually thermalize.
@@
In this Letter, we revisit the question whether non-Abelian gauge theories are chaotic dynamical systems [10, 11, 12], however with a goal to understand the phase space mixing and the process of entropy maximization in an interacting quantum field theory. For this purpose, we measure the maximal Lyapunov exponent of 3 dimensional SU(2) gauge theory using ab-initio lattice techniques in three distinct regimes; for a particular non-thermal classical attractor state [13, 14] and for thermal states near to the color-deconfinement phase transition and at high temperatures, deep in the deconfined phase. Measuring the KS entropy rate in a typical thermal state at $\sim 600$ MeV from the Lyapunov exponents, we for the first time provide an estimate of the typical thermalization time $\sim 0.50(3)$ fm/c for the system to relax into it, starting from the classical attractor regime. This estimate is consistent with the recently reported result [15] for the thermalization time measured in terms of the evolution of the scale that characterizes the infrared magnetic modes of QCD starting from the same non-thermal state. Apart from reinforcing the stochastic features of non-Abelian gauge theories [16] already strongly developed in its classical modes, our study has a practical relevance as well. It provides an estimate of how long it takes to achieve local thermal equilibration i.e. to evolve to a quark gluon plasma phase with initial temperatures $300$-$600$ MeV starting from a highly occupied gluon (glasma) phase believed to be formed in the early stages of a typical heavy-ion collision event [17].
@@ 2 Theoretical background
We begin with a discussion about how to measure the (maximal) Lyapunov exponent in a non-Abelian SU(2) gauge theory which is a non-linear quantum field theory. We recall that in a classical dynamical system, starting from two closely spaced initial coordinates and performing a Hamiltonian evolution if the separation between the two trajectories increases exponentially with time, then the dynamics is defined as chaotic. The rate at which the trajectories diverge from each other is given by the Lyapunov exponent, $\lambda_{L}=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\frac{d(t)}{d(0)}$ where $d(t)$ denotes the separation between two classical trajectories at time $t$. This idea can be easily generalized in a quantum field theory like SU(2), in the semi-classical regime, where the gauge links $U_{\mu}(\mathbf{x})$ are analogous to the space coordinates in a classical configuration space. Using non-perturbative lattice techniques one can generate ensembles of gauge links using Monte-Carlo algorithms. In thermal equilibrium, two different gauge configurations which are separated by infinitesimal Monte-Carlo time steps can be evolved in real-time using the SU(2) Hamiltonian in order to estimate the Lyapunov exponent.
Alternately the Lyapunov exponent can also be motivated by studying the following out-of-time ordered correlation (OTOC) defined in one-dimensional classical dynamics as,
@@
The OTOCs can be easily generalized for quantum field theories. For real-valued scalar fields, the expression in the LHS of Eq. 1 in terms of position and momenta can be replaced by a commutator of the field operator $\hat{\phi}(\mathbf{x},t)$ and their conjugate momentum fields $\hat{\pi}(\mathbf{x},t)$. However for scalar fields the OTOC can be also equivalently defined as $[\hat{\phi}(\mathbf{x},t),\hat{\phi}(\mathbf{0},0)]^{2}$ [18]. In a quantum system, OTOC grows exponentially upto some time $t_{E}$, called the Ehrenfest time [19] beyond which its growth saturates. This is the timescale beyond which the wavefunction spreads over the entire volume of the system and there is a transition from a wave-like behavior to a particle-like behavior of the quantum system [20]. However in the semi-classical limit, the OTOC will not saturate. Such a scenario is realized when scalar fields have large occupation numbers, where one can replace the commutator with the following expression,
@@
An OTOC typically spreads ballistically within the light cone along rays of constant speed $v=x/t$ along the $x$-$t$ plane with a scaling exponent $\nu$ [21] as,
@@
From the spectrum of Lyapunov exponents $\lambda(v)$ in the velocity space of propagating OTOCs, one can define the butterfly velocity $v_{B}$ from the relation $\lambda(v_{B})=0$. The quantities $\lambda_{L},v_{B}$ and exponent $\nu$ are all, in general, temperature dependent. In Section 4 we will extract $\lambda_{L},v_{B}$ for SU(2) gauge theory under thermal equilibrium for a wide range of temperatures as well as in a particular non-equilibrium state. The ab-initio lattice algorithms used for this purpose are detailed in the next section.
@@ 3.1 Studying highly occupied SU(2) gauge theory under non-equilibrium conditions
For a specific realization of a non-equilibrium state of SU(2) gauge theory we start from an initial condition where the momentum $\mathbf{p}$ modes of the gauge fields are occupied according to a phase-space distribution given by, $\tilde{f}(\mathbf{p})=g^{2}f(\mathbf{p})=n_{0}\frac{Q_{s}}{|\mathbf{p}|}\rm{e}^{\frac{-|\mathbf{p}|^{2}}{2Q_{s}^{2}}}$ where $Q_{s}$ is the gluon-saturation scale which is typically between $1$-$2$ GeV and $g$ is the gauge coupling. We consider $Q_{s}=1.5$ GeV in this work. Here and in the subsequent sections we would be considering the theory to be discretized on a 3 dimensional spatial box with $N$ sites along each spatial direction and $a$ is the lattice spacing. The quantity $n_{0}$ sets the initial occupation number of the infrared modes of the gauge fields which we have chosen such that it is close to the typical equilibrium energy densities at high temperatures $T>600$ MeV. Since the occupation numbers are large (owing to g being small), the gauge links behave classically. For generating these classical configurations we solve for the Hamiltonian’s equation of motion for the gauge links $U_{\mathbf{x}}(t)$ and its conjugate fields $E^{ib}_{\mathbf{x}}(t)$ along directions $i=1,2,3$ and with color components $b=1,2$ on the lattice,
@@
where $\Gamma^{b}$ are generators of SU(2) and $V^{ij}_{\mathbf{x}}$ is the spatial staple defined in the $i$-$j$ spatial plane as the derivative of the gauge plaquette $U_{P}$ defined using path ordering $\mathcal{P}$ in color space as,
@@
We use a symplectic leap-frog integrator with a time-step $\Delta t=0.0125Q_{s}$ for the time evolution in order to respect time reversal invariance. We study different lattice volumes with typical extent along the spatial directions $N=128$, with spacings $aQ_{s}=1.115,0.9,0.8$ such that the spatial length of the box in physical units is $\sim 16$ fm. The Gauss law $\frac{a}{\Delta t}\sum_{i}\text{tr}i\Gamma^{a}\left[U_{\mathbf{x}}^{i}(t)-U^{i}_{\mathbf{x}+\hat{i}}(t)\right]$ was implemented with a precession of $10^{-15}$ at $t=0$ and it was checked to remain so at later times. With such a choice of initial condition the system eventually enters a self-similar scaling regime at late times where the gluon distribution function takes the form $\tilde{f}(\mathbf{p},t)=(Q_{s}t)^{-4/7}f_{s}((Q_{s}t)^{-1/7}\mathbf{p})$, where $f_{s}$ is a homogeneous function describing a static distribution [22]. Within such a regime the hard (ultraviolet), electric and the magnetic momentum scales, the later two corresponding to scales set by the Debye mass and the square root of the spatial string tension respectively are well separated similar to a thermal plasma [23].
In order to measure the degree of chaoticity in such a system in a gauge-invariant manner, we follow the procedure as outlined in Ref. [15]. First we implement an infinitesimal shift in the initial condition to $n_{0}+\delta n_{0}$ where $\delta n_{0}=10^{-5}$ and perform a Hamiltonian evolution of the gauge links $U^{\prime}_{\mathbf{x}}(t)$ as a function of time. We next measure the separation between two gauge trajectories at a time $t$ starting with these two infinitesimally close initial conditions through the gauge-invariant distance measure defined in terms of the spatial plaquettes $d(t)=1/N_{P}\sum_{P}|\text{tr}U_{P}(t)-\text{tr}U^{\prime}_{P}(t)|/2$, where $N_{P}$ are the number of independent plaquettes. We particularly choose the time interval to be within the self-similar scaling regime, in order to extract the Lyapunov exponent.
@@ 3.2 Studying thermal SU(2) plasma at high temperatures
At finite temperatures, apart from the hard scale $\pi T$ present in the problem, there are two new scales that are generated dynamically; the electric scale $gT$ and the magnetic scale $g^{2}T/\pi$ [24]. When these three scales are well-separated, one can integrate out the hard and electric gluons, with momenta $\gtrsim gT$ to obtain an effective Hamiltonian [25] which describes the dynamics of the magnetic (soft) gluons where the electric fields $\tilde{E}_{\mathbf{x}}^{i}$ satisfy the equation of motion
@@
where $\sigma$ is the color conductivity and $\tilde{\zeta}^{i}_{\mathbf{x}}(t)$ are the stochastic color-force fields which satisfy $\langle\tilde{\zeta}^{ic}_{\mathbf{x_{1}}}(t_{1})\tilde{\zeta}^{jb}_{\mathbf{x_{2}}}(t_{2})\rangle=2T\sigma\delta^{ij}\delta^{cb}\delta^{3}(\mathbf{x_{1}-x_{2}})\delta(t_{1}-t_{2})$ in accordance to the fluctuation-dissipation relation. Discretizing the Eq. 7, the electric fields, noise fields and color-conductivity can be written in dimensionless units on the lattice in terms of $E_{\mathbf{x}}^{i}=g\tilde{E}_{\mathbf{x}}^{i}a^{2},~{}\zeta^{i}_{\mathbf{x}}=a^{3}g\tilde{\zeta}^{i}_{\mathbf{x}}(t)$ and $\sigma a=\frac{\sigma}{T}.Ta$ respectively. We set $\sigma/T$ to its perturbative values [26] where the coupling $g$ at each temperature $T$ is determined from the 2-loop $\beta$ function. We numerically implement Eq. 7 in dimensionless units which is similar to the electric field evolution in Eq. 5 but with additional contributions due to the noise and damping terms given by $-\Delta t/a\left[a\sigma E^{ib}_{\mathbf{x}}(t-\Delta t)-\mathbf{\zeta}^{ib}_{\mathbf{x}}(t-\Delta t)\right]$. The evolution in Eq. 7 is similar to a Langevin equation of magnetic modes of the gauge fields in presence of a heat bath where the noise term represents the effects of the hard modes which forms the heat-bath and the damping dynamically brings the system to a thermal equilibrium. This classical system consists of $6N^{3}$ oscillators of energy $aT$ and has an energy density on the lattice  [12] given by $\frac{6}{N^{3}}\sum_{\mathbf{k}}|\mathbf{k}|\frac{aT}{|\mathbf{k}|}~{},$ where the sum is over all allowed lattice momenta $\mathbf{k}$, which we have also verified in our calculations. On the other hand, the energy density of quantum SU(2) gauge theory at temperatures $\sim 2$ times than the deconfinement temperature is close to its Stefan-Boltzmann limit $\pi^{2}T^{4}/5$ [27]. The lattice spacing in the effective theory is set in physical units from the criterion that the measured energy density matches with the quantum theory at each $T$. This results in a condition $T.a=(30/\pi^{2})^{1/3}$, which we use to set the lattice spacing in physical units. Once the gauge links are thermalized at a temperature $T$, there is an onset of a plateau in the values of the plaquette term as a function of Langevin time. We then consider two thermal configurations infinitesimally separated as a function of Langevin time, switch off the noise and the damping terms in the Hamiltonian and evolve them subsequently as a function of real-time. We then calculate the gauge invariant distance measure introduced in the earlier section to calculate the (maximal) Lyapunov exponent.

### v2

# Understanding thermalization in a non-Abelian gauge theory in terms of its soft modes

@@ Abstract
We measure the maximal Lyapunov exponent $\lambda_{L}$ of physical states in a SU(2) gauge theory consisting of soft momentum modes both in and out-of-thermal equilibrium conditions using ab-initio lattice techniques. We have implemented different algorithms to appropriately describe the dynamics of soft-modes for a wide range of temperatures and under non-equilibrium conditions. The non-equilibrium state has been realized starting from an over-occupied initial condition for low momentum soft gluons whereas the thermal state comprises of strongly interacting soft gluons at temperatures where these are well separated from the hard momentum modes. Spectra of positive Lyapunov exponents is observed in both these states, similar to a chaotic dynamical system. From the Kolmogorov-Sinai entropy rate measured in terms of this spectrum, we estimate a typical time-scale of $\sim 0.50(3)$ fm/c to achieve thermalization at $T\sim 600$ MeV starting from the non-thermal state. We also measure, for the first time, the $\lambda_{L}$ for long wavelength critical modes of SU(2) using the out-of-time-ordered correlator of a classical $Z_{2}$ scalar field theory, which shares the same universal behavior with SU(2), near the deconfinement phase transition. The $\lambda_{L}$ is observed to maximize at the transition temperature.
@@
Non-equilibrium quantum field theory , QCD , non-Abelian gauge theory , chaos , out-of-time-ordered correlator
@@ 1 Motivation
Understanding the mechanism of thermalization in strongly interacting systems is a challenging research problem in theoretical physics [1, 2] as it would be able to explain physics at different length scales, e.g. in heavy-ion collisions, cold atom traps or in the early universe. Starting from a non-equilibrium state, a necessary condition for eventually reaching a thermal state is a mixing between different regions of the phase space. A mixing in the phase space can occur if there is a mechanical instability in the system [3], which is characterized by at least one positive Lyapunov exponent. Furthermore in the thermal state the entropy or the information content [4] is maximized [5]. Except for a few simple systems [6, 7] it is not very well understood how the entropy evolves as a function of time, starting from a chaotic non-equilibrium state and eventually maximizing itself in a thermal state. One important observable is the rate of change of entropy, known as the Kolmogorov-Sinai (KS) entropy rate [8], which equals the sum of all positive Lyapunov exponents for Hamiltonian systems [9]. This equality is known to be true for dissipative systems that have an attractor solution. Hence calculating the Lyapunov exponents in chaotic dynamical systems is crucial to understand if, and in what time-scales would these eventually thermalize.
@@
In this Letter, we revisit the question whether non-Abelian gauge theories are chaotic dynamical systems [10, 11, 12], however, with a goal to understand the phase space mixing and the process of entropy maximization in an interacting quantum field theory. For this purpose, we measure the maximal Lyapunov exponent of 4-dimensional SU(2) gauge theory using ab-initio lattice techniques in three distinct regimes; for a particular non-thermal classical attractor state [13, 14] and for thermal states near to the color-deconfinement phase transition and at high temperatures, deep in the deconfined phase. Measuring the KS entropy rate of the soft-gluon modes in a typical thermal state at $\sim 600$ MeV from the Lyapunov exponents, we for the first time provide an estimate of the typical thermalization time $\sim 0.50(3)$ fm/c for the system to relax into it, starting from the classical attractor regime. This estimate is consistent with the value that can be obtained using a recently proposed procedure of matching the scales that characterizes the infrared magnetic modes of QCD at equilibrium and in a non-thermal state in the self-similar scaling regime [15]. Apart from reinforcing the stochastic features of non-Abelian gauge theories [16] already strongly developed in its classical modes, our study has a practical relevance as well. Measuring the thermalization time of the chaotic soft-gluon modes thus provides an estimate of how long it takes to achieve local thermal equilibration i.e. to evolve to a quark gluon plasma phase at temperatures $450$-$600$ MeV, a range which is similar to the experimentally measured initial temperature of the fireball formed in a typical heavy-ion collision event [17, 18], starting from a highly occupied gluon phase.
@@ 2 Theoretical background
We begin with a discussion of how to measure the (maximal) Lyapunov exponent in a non-Abelian SU(2) gauge theory which is a non-linear quantum field theory. We recall that, in a classical dynamical system, starting from two closely spaced initial coordinates and performing a Hamiltonian evolution, if the separation between the two trajectories increases exponentially with time then the dynamics is defined as chaotic. The rate at which the trajectories diverge from each other is characterized by the Lyapunov exponent, $\lambda_{L}=\lim_{t\rightarrow\infty}\frac{1}{t}\ln\frac{d(t)}{d(0)}$ where $d(t)$ denotes the separation between two classical trajectories at time $t$. This idea can be easily generalized in a quantum field theory like SU(2), in the semi-classical regime, where the gauge links $U_{\mu}(\mathbf{x})$ are analogous to the space coordinates in a classical configuration space. Using non-perturbative lattice techniques one can generate ensembles of gauge links using Monte-Carlo algorithms. In thermal equilibrium, two different gauge configurations which are separated by infinitesimal Monte-Carlo time steps can be evolved in real-time using the SU(2) Hamiltonian in order to estimate the Lyapunov exponent.
Alternately, another definition of the Lyapunov exponent can also be motivated by studying the following out-of-time ordered correlator (OTOC) defined in one-dimensional classical dynamics as,
@@
The OTOCs can be easily generalized for quantum field theories. For real-valued scalar fields, the expression in the LHS of Eq. 1 can be re-written by replacing the Poisson bracket in terms of position and momenta with a commutator of the field operator $\hat{\phi}(\mathbf{x},t)$ and their conjugate momentum fields $\hat{\pi}(\mathbf{x},t)$. However for scalar fields the OTOC can also be equivalently defined as $[\hat{\phi}(\mathbf{x},t),\hat{\phi}(\mathbf{0},0)]^{2}$ [19]. In a quantum system, OTOC grows exponentially upto some time $t_{E}$, called the Ehrenfest time [20] beyond which its growth saturates. This is the timescale during which the wavefunction has spread over the entire volume of the quantum system and there is a transition from a wave-like to a particle-like behavior [21]. However in the semi-classical limit, the OTOC will not saturate. Such a scenario is realized when scalar fields have large occupation numbers, where one can replace the commutator with the following expression,
@@
An OTOC typically spreads ballistically within the light cone along rays of constant speed $\mathrm{v}=x/t$ in the $x$-$t$ plane with a scaling exponent $\nu$ [22] as,
@@
From the spectrum of the Lyapunov exponents $\lambda(\mathrm{v})$ in the velocity space of propagating OTOCs, one can define the butterfly velocity $\mathrm{v}_{B}$ from the relation $\lambda(\mathrm{v}_{B})=0$. The quantities $\lambda_{L},\mathrm{v}_{B}$ and exponent $\nu$ are all, in general, temperature dependent. The diffusion coefficient $D$ is related to the maximal Lyapunov exponent $\lambda_{L}$ and the butterfly velocity $\mathrm{v}_{B}$ through the relation [23],
$D=\frac{\mathrm{v}_{B}^{2}}{\lambda_{L}}~{}.$ (5)
In Section 4, we will extract $\lambda_{L},\mathrm{v}_{B}$ and $D$ for SU(2) gauge theory in thermal equilibrium for a wide range of temperatures as well as in a particular non-equilibrium state. The ab-initio lattice algorithms used for this purpose are detailed in the next section.
@@ 3.1 Studying highly occupied SU(2) gauge theory under non-equilibrium conditions
For a specific realization of a non-equilibrium state of SU(2) gauge theory, we start from an initial condition where modes of gauge fields the momentum $\mathbf{p}$ are occupied according to a phase-space distribution given by, $\tilde{f}(\mathbf{p})=g^{2}f(\mathbf{p})=n_{0}\frac{Q_{s}}{|\mathbf{p}|}\rm{e}^{\frac{-|\mathbf{p}|^{2}}{2Q_{s}^{2}}}$  [15]. Here $Q_{s}$ is the gluon-saturation scale which is typically between $1$-$2$ GeV and $g$ is the gauge coupling. We consider $Q_{s}=1.5$ GeV in this work. The choice of the initial phase-space distribution becomes irrelevant after a time scale $Q_{s}t\sim n_{0}^{-2}$, which corresponds parametrically to the inverse of the large angle scattering rate between the quasiparticles [24, 25, 14]. Here and in the subsequent sections we would be considering the theory to be discretized on a 3-dimensional spatial box with $N$ sites along each spatial direction and $a$ as the lattice spacing. The quantity $n_{0}$ sets the initial occupation number of the infrared modes of the gauge fields which we have chosen such that it is close to the typical equilibrium energy densities at high temperatures $T\gtrsim 600$ MeV. Since the occupation numbers are large (owing to g being small), the gauge links behave classically. For generating these classical configurations, we solve for the Hamilton’s equations of motion for the gauge links $U^{i}_{\mathbf{x}}(t)$ and its conjugate fields $E^{ib}_{\mathbf{x}}(t)$ on the lattice along directions $i=1,2,3$ and with color components $b=1,2$,
@@
where $\Gamma^{b}$ are generators of SU(2) and $V^{ij}_{\mathbf{x}}$ is the spatial staple defined in the $i$-$j$ plane as the derivative of the gauge plaquette $U_{P}$, defined using path ordering $\mathcal{P}$ in color space as,
@@
The initial conditions for the gauge and electric fields are chosen as a classical-statistical ensemble representing vacuum fluctuations of transversely polarized gluons [24]. We use a symplectic leap-frog integrator for the time evolution in order to respect time-reversal invariance, with a time-step $\Delta t=0.0125Q_{s}$. We study different lattice volumes with extent along the spatial directions $N=128$, with spacings $aQ_{s}=1.115,0.9$ such that corresponding effective temperatures match with the thermal states at temperature $T$ which were generated using the algorithm discussed in the next section i.e. $T_{\mathrm{eff}}\approx T$. The spatial length of the box in physical units is $\sim 16$ fm. The Gauss law $\frac{a}{\Delta t}\sum_{i}\text{tr}i\Gamma^{a}\left[U_{\mathbf{x}}^{i}(t)-U^{i}_{\mathbf{x}+\hat{i}}(t)\right]$ was implemented with a precession of $10^{-15}$ at $t=0$ and it was checked to remain so at later times. With such a choice of initial condition the system eventually enters a self-similar scaling regime at late times where the gluon distribution function takes the form $\tilde{f}(\mathbf{p},t)=(Q_{s}t)^{-4/7}f_{s}((Q_{s}t)^{-1/7}\mathbf{p})$, where $f_{s}$ is a homogeneous function describing a static distribution [14]. Within such a regime the hard, electric and the magnetic momentum scales, the latter two corresponding to scales set by the Debye mass and the square root of the spatial string tension respectively, are well separated similar to a thermal plasma [26].
In order to measure the degree of chaoticity in such a system in a gauge-invariant manner, we follow the procedure as outlined in Ref. [15]. First we implement an infinitesimal shift in the initial condition to $n_{0}+\delta n_{0}$ where $\delta n_{0}=10^{-5}$ and perform a Hamiltonian evolution of the gauge links $U^{\prime}_{\mathbf{x}}(t)$ as a function of time. We next measure the separation between two gauge trajectories at a time $t$ starting with these two infinitesimally close initial conditions, through the gauge-invariant distance measure defined in terms of the spatial plaquettes,
$d(t)=\frac{1}{2N_{p}}\sum_{P}|\text{tr}U_{P}(t)-\text{tr}U^{\prime}_{P}(t)|\sim e^{\lambda_{L}t}$ (8)
where $N_{P}$ are the number of independent plaquettes. We particularly choose the time interval to be within the self-similar scaling regime, in order to extract the Lyapunov exponent.
@@ 3.2 Studying thermal SU(2) plasma at high temperatures
At finite temperatures, apart from the hard scale $\pi T$ present naturally in the problem, there are two new scales that are generated dynamically: the electric scale $gT$ and the magnetic scale $g^{2}T/\pi$ [27]. When these three scales are well-separated, one can integrate out the hard and electric gluons with momenta $\gtrsim gT$ to obtain an effective Hamiltonian [28], which describes the dynamics of the magnetic (soft) gluons with the color electric fields $\tilde{E}_{\mathbf{x}}^{i}$ satisfying the equation of motion
@@
where $\sigma$ is the color conductivity and $\tilde{\zeta}^{i}_{\mathbf{x}}(t)$ are the stochastic color-force fields which satisfy $\langle\tilde{\zeta}^{ic}_{\mathbf{x_{1}}}(t_{1})\tilde{\zeta}^{jb}_{\mathbf{x_{2}}}(t_{2})\rangle=2T\sigma\delta^{ij}\delta^{cb}\delta^{3}(\mathbf{x_{1}-x_{2}})\delta(t_{1}-t_{2})$ in accordance to the fluctuation-dissipation relation. Discretizing the Eq. 9, the electric fields, noise fields and color-conductivity can be written in dimensionless units on the lattice in terms of $E_{\mathbf{x}}^{i}=g\tilde{E}_{\mathbf{x}}^{i}a^{2},~{}\zeta^{i}_{\mathbf{x}}=a^{3}g\tilde{\zeta}^{i}_{\mathbf{x}}(t)$ and $\sigma a=\frac{\sigma}{T}.Ta$ respectively. We set $\sigma/T$ to its perturbative values [29], given by the relation
$\sigma^{-1}=\frac{3N_{c}g^{2}T}{4\pi m_{D}^{2}}\left[\ln\frac{m_{D}}{\gamma}+3.041\right]~{},$ (10)
@@
where the coupling $g$ at each temperature $T$ is determined using the 2-loop $\beta$ function and $1/\gamma$ is the mean free path of the color changing processes which varies between $\gamma/m_{D}\in[0.883,1.034]$ for the range of temperatures considered in our work. Here $N_{c}=2$ and the Debye mass is defined at one-loop order in perturbation theory i.e., $m_{D}=\sqrt{\frac{2}{3}}gT$. We numerically implement Eq. 9 in dimensionless units which is similar to the electric field evolution in Eq. 6 but with additional contributions due to the noise and damping terms given by $-\Delta t/a\left[a\sigma E^{ib}_{\mathbf{x}}(t-\Delta t)-\mathbf{\zeta}^{ib}_{\mathbf{x}}(t-\Delta t)\right]$. The evolution in Eq. 9 is similar to a Langevin equation for magnetic modes of the gauge fields in presence of a heat bath. Here, the noise term represents the effects of the hard modes which form the heat bath and the damping dynamically brings the system to a thermal equilibrium. This classical system consists of $6N^{3}$ oscillators of energy $aT$ and has an energy density on the lattice [12] given by $\frac{6}{N^{3}}\sum_{\mathbf{k}}|\mathbf{k}|\frac{aT}{|\mathbf{k}|}~{},$ where the sum is over all allowed lattice momenta $\mathbf{k}$, which we have also verified in our calculations. On the other hand, energy density of the quantum SU(2) gauge theory at temperatures $\sim 2$ times the deconfinement temperature is close to its Stefan-Boltzmann limit $\pi^{2}T^{4}/5$ [30]. The lattice spacing in the effective theory is set in physical units using the criterion that the measured energy density matches with the quantum theory at each $T$. This results in a condition $T.a=(30/\pi^{2})^{1/3}$, which we use to set the lattice spacing in physical units. Once the gauge links are thermalized at a temperature $T$, there is an onset of a plateau in the values of the plaquette as a function of Langevin time. We then consider two thermal configurations infinitesimally separated as a function of Langevin time, switch off the noise and the damping terms in the Hamiltonian and evolve them subsequently as a function of real-time. This is to implement non-equilibrium evolution of the system of soft gluons along the Schwinger-Keldysh contour, which for this case, consists of only forward-in-time propagation since soft gluons have large occupation numbers. We then calculate the gauge invariant distance measure $d(t)$ introduced earlier to calculate the (maximal) Lyapunov exponent.